# Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature

### v4

@@ Abstract
Photon detection, a fundamental quantum technology, is traditionally treated as a fixed device-level operation governed by intrinsic properties of single-photon detectors (SPDs). High-performance detection has therefore largely relied on superconducting technologies whose requirement for cryogenic operation imposes substantial infrastructure constraints, limiting scalable deployments. Here, the enhanced single-photon detection (ESPD) framework is presented as a system-level theoretical paradigm that shifts photon detection from device-centric optimization to an integrated quantum-information-processing task, by reformulating it as an iteratively enhanced process integrating state-preparation, controlled operations, projective measurements, and multi-copy analysis. ESPD enables systematic performance enhancement through architectural design rather than material modification, thereby circumventing superconducting components, allowing high-performance detection using exclusively room-temperature hardware. Numerical simulations grounded in physically motivated parameters indicate that the ESPD framework can upgrade a conventional room-temperature SPD to effective DE exceeding 93% and DCR below $10^{-9}$, which are comparable to state-of-the-art superconducting SPDs and can significantly relax the minimal tolerable channel transmission rate in quantum communications. While physical realization would require further component integration, this work establishes a rigorous theoretical framework for enhancing detection performance through architectural quantum-information principles, providing a general blueprint for transcending device-level constraints and guiding the development of next-generation room-temperature quantum technologies.
@@
Single-photon detection, System-level quantum architecture, Iterative quantum information processing, Room-temperature quantum technologies, Photonic integration
Hao Shu*
@@
{affiliations}
Dr. Hao Shu
Address: Sun Yat-Sen University, Guangzhou, China
Email Address: Hao_B_Shu@163.com (Primary), shuh6@mail.sysu.edu.cn (Institutional)
@@ 1 Introduction
Single-photon detection plays a central role in quantum optics and serves as the bedrock for a wide range of quantum-information applications. Conventionally, the detection process is treated as a fixed device-level operation primarily determined by the material properties of single-photon detectors (SPDs). The primary performance metrics characterizing an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former directly dictates the success probability of quantum operations conditioned on detection events, while the latter imposes a fundamental signal-to-noise constraint in high-attenuation scenarios.
@@
In terms of DE, high-performance detection is indispensable for retrieving meaningful outcomes in quantum experiments. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the overall success probability decreases exponentially with imperfect DE [1, 2, 3]. In fundamental tests of quantum nonlocality, verifying violations of the Clauser-Horne-Shimony-Holt or Eberhard inequalities requires DE exceeding strict thresholds to close detection loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where one photon of an entangled pair signals the presence of the other, the achievable heralding rate is also strictly limited by DE [13, 14, 15, 16].
@@
With respect to DCR, a particularly stringent constraint arises in quantum communication, notably in quantum key distribution (QKD). The security of practical QKD protocols relies on maintaining the quantum bit error rate (QBER) below protocol-dependent thresholds that are significantly lower than 50% [17, 18, 19, 20]. However, optical communication channels inevitably induce exponential photon attenuation with distance [21, 22, 23, 24]. When the effective signal rate, determined by channel transmission for a given DE, approaches the DCR floor of the SPDs, detection events become dominated by dark counts, which contribute a 50% error rate, causing the QBER to rise precipitously. Consequently, the achievable secure distance of QKD systems is ultimately bounded by the DCR of the employed SPDs [25, 26, 27].
@@
Motivated by these critical demands, the development of high-performance SPDs has remained a primary research focus. While conventional semiconductor-based SPDs offer the advantage of non-cryonic operation, they typically suffer from an intrinsic trade-off between limited DE and relatively high DCR. For instance, a frequency upconversion SPD demonstrated in 2004 operated at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP SPD reported in 2017 operated at 223 K achieved a DE of about 27.5% with a DCR of $10^{-6}$ [29]. Such specifications often fall short of the requirements for advanced quantum protocols, which frequently demand DE exceeding 90% and DCR below $10^{-7}$.
@@
To meet these stringent requirements, substantial effort has shifted toward designing SPDs with superconducting technologies, most notably superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These devices have successfully pushed DE beyond 90% while suppressing DCR by several orders of magnitude compared to semiconductor counterparts, thereby enabling demanding applications such as long-distance QKD [19, 34, 35, 36]. Recent representative benchmarks of state-of-the-art superconducting SPDs are summarized in Table 1.
Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
Year Researchers (et al.) Operation Temperature Detection Efficiency Dark Count Rate
@@
To address the limitations of room-temperature SPDs without resorting to cryogenic conditions, the enhanced single-photon detection (ESPD) framework is introduced, shifting the focus from device-centric optimization to a system-level architectural paradigm. Within this framework, single-photon detection is reformulated as a composite quantum-information-processing task that integrates state preparation, controlled quantum operations, projective measurements, and multi-copy decision analysis, rather than relying solely on the material properties of a single SPD. Through structured system-level integration, the ESPD framework enables effective enhancement of detection performance using exclusively room-temperature components.
Numerical simulations, grounded in physically motivated parameters drawn from commercially available devices and experimentally reported SPDs, indicate that, under an exclusively room-temperature component setting, the ESPD framework enables effective DE exceeding 93% together with effective DCR below $10^{-9}$, starting from a legacy SPD (for example, the one reported two decades ago with about 59% DE and $10^{-2}$ DCR [28]). These effective performance metrics are comparable to those of state-of-the-art superconducting SPDs and lead to a substantial relaxation of the minimal tolerable channel transmission rate in QKD.
While physical implementation would require further system-level integration of quantum technologies, the ESPD framework provides a structured theoretical basis for achieving high-performance photon detection at room temperature. By explicitly analyzing technological feasibility and the trade-off between resource overhead and operating temperature, the ESPD framework offers a feasible technological roadmap and practical guidance for future experimental efforts. It serves as a theoretical blueprint for overcoming intrinsic device limitations, with the potential to unlock a wide range of quantum applications in environmentally constrained scenarios.
2 Method and Result
This section details the architectural structure of the ESPD framework and formulates the corresponding detection process. The ESPD paradigm is introduced in Section 2.1, with theoretical formulations established in Section 2.2. Analytical approximations are developed in Section 2.3, and numerical simulations are presented in Section 2.4. Finally, implications for QKD are discussed in Section 2.5.
@@ 2.1 Paradigm of the Enhanced Single-Photon Detection
The ESPD architecture is formulated as a level-by-level framework in which the effective detection performance of a given SPD is enhanced through successive processing stages. Here, the enhancement is quantified in terms of DE and DCR, defined respectively as the probability of generating a positive detection event given a non-vacuum (single-photon) input and the probability of a positive report given a vacuum input.
Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with effective DE $\eta_{s}$ and effective DCR $d_{s}$. The iteration begins with an existing SPD, denoted as $\text{ESPD}_{0}$, characterized by the initial values $\eta_{0}=\eta$ and $d_{0}=d$.
With qubit logic, the schematic of the ESPD paradigm is illustrated in Fig. 1.
Figure 1: Schematic representation of the recursive ESPD architecture. The $(s+1)$-th level detector ($\text{ESPD}_{s+1}$) is constructed by integrating multiple copies of the $s$-th level detector ($\text{ESPD}_{s}$) with controlled gates. Here, $C(F_{i})$ are controlled gates, SP are state-preparation operations, measurements are projective measurements via basis $\{|0\rangle,|1\rangle\}$, and filters represent that the $|0\rangle$ path is aborted.
The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, is defined as follows. As a representative example, an implementation may be considered in which the degree of freedom (DOF) is photon polarization and the controlled operations correspond to controlled-NOT (C-NOT) gates.
@@
    DOF Selection: A manipulable DOF of the incoming photon is chosen (for example, polarization). The computational basis for this DOF is denoted by $\{|0\rangle,|1\rangle\}$. The signal is modeled as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle$ and $|1\rangle$ denote orthonormal states of a non-vacuum photon [^1] If the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents a non-vacuum state, reducing the description to qubit. In practical architectures, an internal DOF such as polarization, distinct from photon existence, is often more convenient..
@@
    State Preparation: A state-preparation operation is applied to initialize the selected DOF to $|1\rangle$. As a result, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum input remains in state $|\emptyset\rangle$.
@@
    Controlled Operations: A set of $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates), are applied to couple the signal path with $n_{s+1}$ auxiliary photons, each initialized in state $|0\rangle$.
@@
    Auxiliary Measurement: Projective measurements are performed on the auxiliary outputs in basis $\{|0\rangle,|1\rangle\}$. Outcome paths corresponding to $|0\rangle$ are discarded, and only the $|1\rangle$ outcomes are retained for subsequent detection.
@@
    Detection: Detection is carried out on all remaining paths, including both the signal path and the auxiliary paths, using $s$-th level detectors ($\text{ESPD}_{s}$).
@@
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detection events are registered among the $n_{s+1}+1$ detectors, consisting of $n_{s+1}$ auxiliary detectors and one detector on the signal path. Otherwise, a negative report is returned. Here, $k_{s+1}$ is a preselected threshold integer, and $n_{s+1}$ denotes the number of effective controlled operations [^2] The main discussions of the article assume deterministic controlled gates, for which $n_{s+1}$ equals the number of implemented operations. For probabilistic gates, $n_{s+1}$ refers to the number of successful operations after post-selection. In this case, the decision logic is naturally reformulated in terms of the fraction of positive detection outcomes among the effective detections exceeding a predefined threshold..
Physical intuition: From a detection perspective, each level of the ESPD framework increases the number of effective detection opportunities available to an input signal. In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level produces $n_{s+1}$ effective signal instances through the controlled modules, which are independently detected by $n_{s+1}$ SPDs. A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered. Under this idealized assumption, the effective DE increases with decreasing threshold $k_{s+1}$; for instance, when $k_{s+1}=1$, the effective DE is given by $1-(1-\eta_{s})^{n_{s+1}}$. Furthermore, the effective DCR is suppressed, since a false positive outcome requires dark-count events to occur in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$. This physical intuition extends beyond the idealized limit when imperfections in the auxiliary components introduce errors that remain small compared to the activation probability of a genuine input signal. A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section 2.2.
@@ 2.2 Theoretical Formulations
The rigorous theoretical formulation of the ESPD framework is developed in this subsection. The relevant notations and assumptions are introduced first, followed by the exact analytical expressions and the approximation analysis.

### v5

# From Device to Dynamics: An Iterative Architectural Framework for High-Performance Single-Photon Detection at Room Temperature

@@ Abstract
Photon detection is a cornerstone of quantum technology, traditionally regarded as a static device-level operation constrained by the intrinsic physical properties of single-photon detectors (SPDs). Consequently, high-performance detection has been heavily reliant on superconducting technologies, whose requirement for cryogenic temperatures imposes significant infrastructure burdens and limits scalable deployment. To circumvent these constraints, we propose the Enhanced Single-Photon Detection (ESPD) framework, which shifts the photon-detection paradigm from device-centric optimization to an integrated quantum-information-processing (QIP) task. By incorporating state preparation, controlled operations, projective measurements, and multi-copy decision analysis, we establish a nonlinear dynamical model that reformulates detection as an iteratively enhanced process. This architecture enables systematic performance upgrades through structural design rather than material modification, allowing high-performance detection with exclusively room-temperature hardware. Through analytical approximations, Monte Carlo analysis, and numerical simulations, grounded in parameters derived from non-superconducting components, we show that the ESPD dynamics converge to a high-performance basin of attraction even when initialized by low-performance SPDs. Specifically, the framework can upgrade a conventional room-temperature SPD to achieve an effective detection efficiency (DE) exceeding 93% with a dark count rate (DCR) below $10^{-9}$, metrics comparable to state-of-the-art superconducting nanowire SPDs (SNSPDs). Such enhancements significantly lower the threshold for the required minimum channel transmission rate in quantum communication applications. While physical realization requires further component integration efforts, this work establishes a rigorous theoretical foundation for enhancing detection via architectural QIP principles. It provides not only a blueprint for next-generation room-temperature photon detection but also a general methodology for transcending device-level constraints in broader quantum technologies.
@@
Single-photon detection, Room-temperature quantum technologies, Iterative enhanced quantum information processing, Nonlinear dynamics, Photonic integration
@@
organization=Sun-Yat-Sen University,city=Shenzhen, country=China
@@ 1 Introduction
Photon detection is a foundational pillar of quantum optics, providing essential interfaces for nearly all quantum information applications. Traditionally, the detection process has been conceptualized as a static, device-level operation, whose performance is regarded as intrinsic to the single-photon detector (SPD) material properties. The utility of an SPD is primarily governed by two competing metrics: detection efficiency (DE, denoted by $\eta$), which represents the probability of detecting a non-vacuum incident photon, and the dark count rate (DCR, denoted by $d$), which represents the probability of returning an incorrect positive detection report in the absence of a photon.
@@
High DE is indispensable for obtaining meaningful outcomes in quantum experiments. In fundamental tests of quantum nonlocality, verifying violations of the Clauser-Horne-Shimony-Holt or Eberhard inequalities requires DE exceeding strict thresholds to close detection loopholes [1, 2, 3, 4, 5, 6, 7], such as approximately 67% [8] or 83% [9]. In optical quantum computation, where successful events can be probabilistic and require the simultaneous detection of multiple photons, the overall success probability decreases exponentially with imperfect DE [10, 11, 12]. In heralded single-photon sources, where one photon of an entangled pair signals the presence of the other, the achievable heralding rate is also strictly limited by DE [13, 14, 15, 16].
@@
Regarding DCR, a particularly stringent constraint arises in quantum communication, notably quantum key distribution (QKD). The security of practical QKD protocols relies on maintaining the quantum bit error rate (QBER) below protocol-dependent thresholds that are usually lower than 11% [17, 18, 19, 20, 21]. However, optical communication channels inevitably induce exponential photon attenuation with distance [22, 23, 24, 25]. When the effective signal rate, determined by exponentially decay channel transmission for a given DE, approaches the DCR floor of the SPDs, detection events become dominated by dark counts that contribute a 50% error rate, causing the QBER to rise precipitously. Consequently, the achievable secure distance of QKD systems is ultimately bounded by the DCR of the employed SPDs [18, 26, 27].
@@
Motivated by these critical demands, the development of high-performance SPDs has remained a primary research focus. While conventional semiconductor-based SPDs offer the advantage of non-cryogenic operation, they typically suffer from intrinsic limits on relatively low DE and high DCR. For instance, a frequency upconversion SPD reported in 2004 operates at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP SPD reported in 2017 operates at 223 K achieved a DE of about 27.5% with a DCR of $10^{-6}$ [29]. Such specifications often fall short of advanced quantum task requirements, which frequently demand DE exceeding 90% and DCR below $10^{-7}$.
@@
To meet these stringent requirements, substantial efforts have gravitated toward superconducting SPD technologies, most notably superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These devices have successfully pushed DE beyond 90% while suppressing DCR by several orders of magnitude compared to semiconductor counterparts, thereby enabling demanding applications such as long-distance QKD [20, 34, 35, 36]. Recent representative benchmarks of state-of-the-art superconducting SPDs are summarized in Table 1.
Table 1: Recent representative benchmarks of superconducting SPDs. OT: Operating Temperature; DE: Detection Efficiency; DCR: Dark Count Rate
Year Researchers (et al.) OT DE DCR
@@
To address the limitations of room-temperature SPDs without resorting to cryogenic conditions, we introduce the enhanced single-photon detection (ESPD) framework, which shifts the focus from device-centric optimization to a system-level architectural paradigm with nonlinear dynamics. Within this framework, photon detection is reformulated as a composite quantum-information-processing task that integrates state preparation, controlled quantum operations, projective measurements, and multi-copy decision analysis, rather than relying solely on SPD material properties. By formulating this multi-stage enhancement approach as a non-linear dynamical system, we analyze the convergence range and steady-state performance of the ESPD structure, providing a novel perspective on hardware noise suppression through iterative convergence processing. Our analysis demonstrates that this framework can elevate the performance of conventional, room-temperature SPDs to levels comparable to, or even exceeding, those of cryogenic superconducting counterparts.
Numerical simulations, grounded in physically motivated parameters from commercially available devices, indicate that the iterative mapping within the ESPD framework exhibits strong convergence toward high-performance regimes. Monte Carlo analysis reveals that, provided the initial SPD parameters fall within a broad basin of attraction (i.e., not starting with an extremely poor SPD), the evolved DE can surpass 90% alongside ultra-low DCR. For instance, starting from a legacy SPD (e.g., DE $\approx 59\%$ and DCR $\approx 10^{-2}$ [28]), the system converges to an effective DE exceeding 93% and an effective DCR below $10^{-9}$. Such performance rivals or even outperforms state-of-the-art superconducting SPDs, substantially relaxing the minimal tolerable channel transmission rate in QKD, and is achieved under room-temperature settings, circumventing the need for cryogenic infrastructure.
While physical implementation will require further system-level integration of quantum technologies, the ESPD framework provides a structured nonlinear dynamical formulation for room-temperature high-performance photon detection. Furthermore, we offer a practical roadmap to guide future experimental efforts by explicitly discussing technological feasibility and analyzing the trade-offs between resource overhead and operating temperature. Ultimately, the design principles and perspective shift offered by the ESPD framework serve as a theoretical blueprint for overcoming intrinsic device limitations through dynamical process design, which can boost the development of a broader range of quantum devices beyond photon detection and unlock diverse high-performance quantum applications at room temperature.
2 Methods and Results
This section details the ESPD architecture and formulates the corresponding detection dynamics. The ESPD paradigm is introduced in Section 2.1, with rigorous theoretical formulations established in Section 2.2. Analytical approximations are developed in Section 2.3, followed by (Monte Carlo) stability analysis and numerical simulations in Section 2.4. Finally, the implications for QKD applications are discussed in Section 2.5.
@@ 2.1 Paradigm of the Enhanced Single-Photon Detection
The ESPD architecture is formulated as a recursive, level-by-level framework in which baseline SPD performance is systematically enhanced through successive processing stages. Within this paradigm, the performance is characterized by the effective DE ($\eta_{s}$) and DCR ($d_{s}$) at level $s$. Recall that DE is the probability of generating a positive detection event given a non-vacuum (single-photon) input, and DCR is the probability of generating an (incorrect) positive report given a vacuum input. Let $\text{ESPD}_{s}$ denote the established detector at level $s$, characterized by $(\eta_{s},d_{s})$. The evolution begins with an existing SPD, denoted as $\text{ESPD}_{0}$, with initial values $\eta_{0}=\eta$ and $d_{0}=d$, the performance of the physical baseline SPD. The schematic of this recursive logic is illustrated in Fig. 1.
Figure 1: Schematic representation of recursive ESPD architecture. The $(s+1)$-th level detector ($\text{ESPD}_{s+1}$) is constructed by integrating multiple copies of the $s$-th level detector ($\text{ESPD}_{s}$) with state preparation, controlled gates, and projective measurements. Here, $C(F_{i})$ are controlled gates, SP are state-preparation operations, measurements are projective in the basis $\{|0\rangle,|1\rangle\}$, and filters indicate the abortion of the $|0\rangle$ path.
The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, is defined as follows. As a representative example, an implementation can be considered with the degree of freedom (DOF) using photon polarization and the controlled operations using controlled-NOT (C-NOT) gates.
@@
    DOF Selection: Select a manipulable DOF of the incoming photon (for example, polarization). The computational basis for this DOF is denoted by $\{|0\rangle,|1\rangle\}$. Hence, the signal is modeled as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle$ and $|1\rangle$ denote orthonormal states of a non-vacuum photon [^1] If the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents a non-vacuum state, reducing the description to qubit. In practical architectures, an internal DOF such as polarization, distinct from photon existence, is often more convenient..
@@
    State Preparation: Apply the state-preparation operation, initializing the selected DOF to $|1\rangle$. As a result, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum input remains in state $|\emptyset\rangle$.
@@
    Controlled Operations: Operate a set of $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates), coupling the signal path with $n_{s+1}$ auxiliary photons, each initialized in state $|0\rangle$.
@@
    Auxiliary Measurement: Perform projective measurements on the auxiliary outputs in basis $\{|0\rangle,|1\rangle\}$. Outcome paths corresponding to $|0\rangle$ are discarded; only the $|1\rangle$ outcomes are retained for subsequent detections.
@@
    Detection: Carried out detections on all remaining paths, including both the signal path and the auxiliary paths, using $s$-th level detectors ($\text{ESPD}_{s}$).
@@
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detection events occur among the $n_{s+1}+1$ detectors, consisting of $n_{s+1}$ auxiliary detectors and one detector on the signal path. Otherwise, a negative report is returned. Here, $k_{s+1}$ is a preselected threshold integer, and $n_{s+1}$ denotes the number of effective controlled operations [^2] The main discussions of the article assume deterministic controlled gates, for which $n_{s+1}$ equals the number of implemented operations. For probabilistic gates, $n_{s+1}$ refers to the number of successful operations after post-selection. In this case, the decision logic is naturally reformulated in terms of the fraction of positive detection outcomes among the effective detections exceeding a predefined threshold..
Physical intuition: From a functional perspective, each level of the ESPD framework increases the number of effective detection opportunities available to an input signal. In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level produces $n_{s+1}$ auxiliary signal instances identical to the input one through the controlled modules, which are independently detected by the $n_{s+1}$ ESPDs. A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered among the $n_{s+1}+1$ detections. Therefore, the effective DE increases with decreasing threshold $k_{s+1}$; for instance, when $k_{s+1}=1$, the effective DE is given by $1-(1-\eta_{s})^{n_{s+1}}$. Furthermore, the effective DCR is suppressed because a false positive outcome requires at least $k_{s+1}$ independent dark count events, leading to an approximate probability of $d_{s}^{k_{s+1}}$. This architectural logic transforms detection from a single-shot device event into a more robust collective process, where multi-copy analysis can suppress the errors introduced by SPD imperfections. This intuition extends beyond the idealized limit when imperfections in the auxiliary components introduce errors that remain small compared to the activation probability of a genuine input signal. A rigorous treatment of these dynamics, accounting for non-ideal devices and error propagation, follows in Section 2.2.
@@ 2.2 Theoretical Formulations
This subsection develops the rigorous theoretical foundation of the ESPD framework. We first define the governing parameters and underlying assumptions, followed by deriving the analytical expressions of the ESPD dynamics.